# Internal and String Stability of an Observer-based Controller for Vehicle Platooning under the MPF Topology

### v2

@@ VI-B2 Condition (32) is not satisfied
The example in Sec. VI-B shows that given a fixed time headway $h$, there exist multiple solutions of $\alpha$ and $b$ to guarantee the string stability of vehicle platooning. Sec. VI-A demonstrates the proposed observer-based controller controlling performance is better than the PID controller in [6]. Note that only two parameters, namely $\alpha$ and $b$, need to be tuned with the help of the feasible region, compared to at most three parameters in PID controller.
VI-C Algorithm 1 for controller parameters with time headway minimization in Sec. V-D
@@
One can observe that values of $\alpha$ in the above two examples in Secs. VI-B and VI-C are different. The reason is that Algorithm 1 verified in Sec. VI-C is by fixing $\alpha=2\tau$ to minimize the value of $h$. Actually, given a fixed $h$, $\alpha$ can have different values; and this is the heuristic searching algorithm in Sec. VI-B demonstrates. To sum up, the objective of the heuristic searching algorithm is to design controller parameters $\alpha$ and $b$ given a fixed $h$ while the objective of Algorithm 1 is to minimize the value of $h$ by fixing $\alpha=2\tau$. To analyze the relationship among $\alpha,b$ and $h$ directly without fixing one value would be an interesting direction in the future.
VI-D Performance of observer-based controller related to different values of $h$
@@
VII Conclusions and Future Directions
VII-A Conclusions
@@ VII-B Future Directions
    As it is stated in [40], string stability is only a necessary but not a sufficient condition for collision avoidance; studying the platoon’s collision avoidance conditions is of practical importance.

### v3

@@ VII-B2 Condition (32) is not satisfied
The example in Sec. VII-B shows that given a fixed time headway $h$, there exist multiple solutions of $\alpha$ and $b$ to guarantee the string stability of vehicle platooning. Sec. VII-A demonstrates the proposed observer-based controller controlling performance is better than the PID controller in [6]. Note that only two parameters, namely $\alpha$ and $b$, need to be tuned with the help of the feasible region, compared to at most three parameters in PID controller.
VII-C Algorithm 1 for controller parameters with time headway minimization in Sec. V-D
@@
One can observe that values of $\alpha$ in the above two examples in Secs. VII-B and VII-C are different. The reason is that Algorithm 1 verified in Sec. VII-C is by fixing $\alpha=2\tau$ to minimize the value of $h$. Actually, given a fixed $h$, $\alpha$ can have different values; and this is the heuristic searching algorithm in Sec. VII-B demonstrates. To sum up, the objective of the heuristic searching algorithm is to design controller parameters $\alpha$ and $b$ given a fixed $h$ while the objective of Algorithm 1 is to minimize the value of $h$ by fixing $\alpha=2\tau$. To analyze the relationship among $\alpha,b$ and $h$ directly without fixing one value would be an interesting direction in the future.
VII-D Controller performance related to different values of $h$
@@
VIII Conclusions and Future Directions
VIII-A Conclusions
@@ VIII-B Future Directions
    As it is stated in [40], string stability is only a necessary but not a sufficient condition for collision avoidance/safety; studying further the platoon’s collision avoidance/safety conditions is of practical importance.
@@ Appendix C The details of steps i and ii of proposed heuristic searching algorithm
ACKNOWLEDGMENT
The authors would like to thank anonymous reviewers and the Associate Editor for their important, enlightening and valuable comments.